# Understanding the Local Geometry of Generative Model Manifolds

### v1

@@ Appendix C Extra Figures
Figure 19: After 11395 optimization steps. Geometry of a diffusion model input-output mapping, trained to on a toy 2D distribution. Local scaling lower around data manifold, local complexity higher around manifold, rank is lower around manifold as well. t=50 has considerably low variance in local scaling showing that final timestep has a diminishing change of density.
@@
Figure 20: Reward guidance on stable diffusion (maximizing the reward).
@@
$\psi$
$\delta$
$\nu$
Figure 27: Vendi score [5] calculated for samples from different local descriptor level sets of a Beta-VAE. We take upto $150$ samples from each level set and compute vendi score seperately for the MNIST train dataset, test dataset and generated samples.
$\psi$
$\delta$
$\nu$
Figure 28: Sub-population differences of local descriptors in training data. Class-wise local descriptor distributions for a Beta-VAE trained unconditionally on MNIST. We use the VAE encoder to obtain latent vectors for the training dataset and compute descriptors using the VAE decoder. We can observe class specific separation in the distributions indicating that geometric properties may be different for different sub-populations within the training dataset.

### v2

@@ Appendix G Extra Figures
Figure 21: 2D Data Manifold Geometry, A toy Example. After 11395 optimization steps. Geometry of a diffusion model input-output mapping, trained to on a toy 2D distribution. Local scaling lower around data manifold, local complexity higher around manifold, rank is lower around manifold as well. t=50 has considerably low variance in local scaling showing that final timestep has a diminishing change of density.
@@
Figure 22: Reward guidance on stable diffusion (maximizing the reward).We observe a significant increase in both background detail and artifact diversity within the generated images.
a Australian terrier with
a city in the background
$\rho=0$                    $\rho=1$                    $\rho=2$                    $\rho=3$
a Australian terrier with a city
in the background
$\rho=0$                    $\rho=-1$                    $\rho=-3$                    $\rho=-5$
Figure 23: Controlling mage diversity with local scaling. Using Reward guidance to increase (top row) and decrease diversity (bottom row) using same initial seed.
a Australian terrier
in the snow
$\rho=0$                    $\rho=1$                    $\rho=2$                    $\rho=3$
a Australian terrier
in the snow
$\rho=0$                    $\rho=-1$                    $\rho=-3$                    $\rho=-5$
Figure 24: Controlling mage diversity with local scaling.Using Reward guidance to increase (top row) and decrease diversity (bottom row) using same initial seed.
@@
Figure 31: UMAP visualization of the aggregated local geometry descriptors (local smoothness, local rank, and local complexity). This reveals distinct, non-overlapping clusters, clearly separating the Imagenet, Imagenet Corrupted with Gaussian Noise, and Chest X-ray datasets. This visual evidence underscores the effectiveness of aggregating the descriptors to capture unique patterns within each dataset, demonstrating its ability to provide a meaningful and interpretable representation of the underlying data
Training Samples
Generated Samples
Local scaling, $\psi_{t=.18T}$
Figure 32: Highest local scaling level sets are anti-modes. In all of our experiments, we have observed that samples from the highest local scaling level sets have lower vendi score. We repeat the experiments from Fig. 14 for a mixture of nine gaussians. We see that the regions between the gaussian modes inside the domain of the data, i.e., the anti-modes, have higher local scaling, with the highest local scaling values at the center of the anti-modes.